### v1

# Adaptive Monitoring of Stochastic Fire Front Processes via Information-seeking Predictive Control

Abstract
We consider the problem of adaptively monitoring a wildfire front using a mobile agent (e.g., a drone), whose trajectory determines where sensor data is collected and thus influences the accuracy of fire propagation estimation. This is a challenging problem, as the stochastic nature of wildfire evolution requires the seamless integration of sensing, estimation, and control, often treated separately in existing methods. State-of-the-art methods either impose linear-Gaussian assumptions to establish optimality or rely on approximations and heuristics, often without providing explicit performance guarantees. To address these limitations, we formulate the fire front monitoring task as a stochastic optimal control problem that integrates sensing, estimation, and control. We derive an optimal recursive Bayesian estimator for a class of stochastic nonlinear elliptical-growth fire front models. Subsequently, we transform the resulting nonlinear stochastic control problem into a finite-horizon Markov decision process and design an information-seeking predictive control law obtained via a lower confidence bound-based adaptive search algorithm with asymptotic convergence to the optimal policy.
I Introduction
The 2025 Southern California wildfires were among the most devastating in the state’s history: the Palisades Fire in Los Angeles and the Eaton Fire in Altadena burned nearly 40,000 acres, destroyed over 16,000 structures, and displaced hundreds of thousands of people [7]. Accurate estimation of wildfire propagation is therefore critical for effective disaster response and informed decision-making. Motivated by this need, this work investigates adaptive monitoring of a stochastic wildfire front using a mobile agent.
The task requires planning the agent’s trajectory over a rolling finite horizon to minimize uncertainty in estimating the fire’s evolution from sensor data. At each step, the agent re-plans based on the current environment, yielding a complex problem that demands an integrated approach to sensing, estimation, and control. Existing methods often address only parts of this problem, either by decoupling sensing, estimation, and control, or by simplifying assumptions [11, 15]. The problem considered in this work relates to active sensing [9], information gathering [1], and sensor management [5]. Sensor management approaches typically address stateless sensors without dynamics, focusing on placement or node selection, and are therefore limited in scenarios where sensor states evolve (e.g., drones equipped with onboard sensors). In such dynamic settings, adaptive control strategies are required. For stochastic dynamic processes, many methods adopt myopic control [3], whereas non-myopic schemes are typically greedy [13] or heuristic with sub-optimality guarantees [6]. Approaches that provide optimality guarantees often assume linear dynamics with Gaussian noise and apply the certainty equivalence principle (CEP), thereby reducing the problem to deterministic optimal control [8]. These assumptions, however, do not hold for the problem addressed in this work.
In summary, this work integrates sensing, estimation, and control within a unified stochastic optimal control (SOC) framework for adaptive wildfire-front monitoring using a mobile agent. We develop a recursive Bayesian estimator for elliptical fire-front dynamics under limited sensing and uncertainty, and reformulate the nonlinear SOC problem as a finite-horizon Markov decision process (MDP). The MDP is solved via a lower-confidence-bound (LCB) guided adaptive search that asymptotically converges to the optimal policy.
II Problem Formulation
II-A Problem Objective
Let a mobile agent be at state $y_{t}$ at time step $t$, and let the agent’s belief distribution over the fire front state $X_{t}$ be denoted by $\mathcal{B}_{t}(X_{t}|Z_{1:t})$ (abbreviated as $\mathcal{B}_{t}$) which was computed using measurements $Z_{1:t}=[Z_{1},\ldots,Z_{t}]$ up to time step $t$. The objective is to compute the optimal sequence of control inputs $\{u_{t|t},\ldots,u_{t+T-1|t}\}$ over a finite rolling planning horizon of $T$ time steps that minimizes:
$\mathbb{E}_{Z_{t}^{1:T}}\left\{\sum_{\tau=1}^{T}\mathcal{C}_{\tau}\left(\mathcal{B}^{-}_{t+\tau|t},Z_{t+\tau|t}(u_{t+\tau-1|t})\right)\right\}.$ (1)
Here, the subscript $t+\tau|t$ denotes predicted quantities at time step $t+\tau$ for $\tau\in\{1,\dots,T\}$ within the planning horizon, based on information available at time step $t$. The cost function $\mathcal{C}_{\tau}$ is a bounded, real-valued function that takes as input: a) the predictive density $\mathcal{B}^{-}_{t+\tau|t}$ i.e., $\mathcal{B}^{-}_{t+\tau|t}(X_{t+\tau|t}|Z_{1:t+\tau-1|t})$, of the fire front state at time $t+\tau|t$, and b) the future i.e., predicted, measurement $Z_{t+\tau|t}(u_{t+\tau-1|t})$, which will be received given that the agent executes the control input $u_{t+\tau-1|t}$ and moves to the predicted state $y_{t+\tau|t}$. It then returns a value representing the uncertainty of the fire front state captured in the resulting (pseudo) posterior belief $\mathcal{B}_{t+\tau|t}$. An information-rich measurement set is one that reduces the dispersion in the posterior, which is the agent’s objective over the planning horizon. The expectation is taken with respect to the future measurement set $Z_{t}^{1:T}=\{Z_{t+1|t},\ldots,Z_{t+T|t}\}$. Subsequently, the agent executes the first control input in the sequence, i.e., $u_{t|t}$, transitions to its new state $y_{t+1}$, receives the real measurement $Z_{t+1}$, computes the posterior belief $\mathcal{B}_{t+1}(X_{t+1}|Z_{1:t+1})$, and repeats the process described above for time step $t+1$.
II-B Fire Front Propagation Model
In this work, as we discuss next, we employ a stochastic adaptation of the deterministic elliptical fire propagation model proposed in [12]. This model, which is currently utilized in various fire-area simulators [4], describes the spatiotemporal evolution of a fire front using a nonlinear system of first-order differential equations. Specifically, the fire front is represented as an ellipse defined by a series of $N$ vertices that collectively delineate the propagating fire’s edge at a specific moment. The spatiotemporal discrete-time dynamics of vertex $i\in\{1,\ldots,N\}$ at time step $t$ are given by:
$x^{i}_{t}=x^{i}_{t-1}+\Delta t\,\dot{x}^{i}_{t-1},$ (2)
where $x^{i}_{t}=[\text{x}^{i}_{t},\text{y}^{i}_{t}]^{\top}\in\mathbb{R}^{2}$ is the state of vertex $i$ composed of 2D Cartesian coordinates, $\Delta t$ is the sampling interval, and the fire growth velocity at vertex $i$ is given by $\dot{x}^{i}_{t-1}=$
$\begin{bmatrix}\displaystyle\frac{\alpha_{1}^{2}(i)\cos\left(\theta(i)\right)SC(i)-\alpha_{2}^{2}(i)\sin\left(\theta(i)\right)CS(i)}{\sqrt{\alpha_{2}^{2}(i)CS(i)^{2}+\alpha_{1}^{2}(i)SC(i)^{2}}}+C_{1}(i)\!\\ \displaystyle\frac{-\alpha_{1}^{2}(i)\sin\left(\theta(i)\right)SC(i)-\alpha_{2}^{2}(i)\cos\left(\theta(i)\right)CS(i)}{\sqrt{\alpha_{2}^{2}(i)CS(i)^{2}+\alpha_{1}^{2}(i)SC(i)^{2}}}+C_{2}(i)\!\end{bmatrix}$,
where $SC(i)=\text{x}^{i}_{s}\sin\left(\theta(i)\right)+\text{y}^{i}_{s}\cos\left(\theta(i)\right)$, $CS(i)=\text{x}^{i}_{s}\cos\left(\theta(i)\right)-\text{y}^{i}_{s}\sin\left(\theta(i)\right)$, $C_{1}(i)=\alpha_{3}(i)\sin\left(\theta(i)\right)$, $C_{2}(i)=\alpha_{3}(i)\cos\left(\theta(i)\right)$, and $\text{x}^{i}_{s},\text{y}^{i}_{s}]^{\top}$ are the components of the tangent vector at vertex $i$, providing the local orientation of the fire front at that point.
Environmental conditions, such as fuel type and weather, local to each vertex, affect the forward fire propagation rate and direction. These factors include wind direction and speed, denoted by $\theta$ and $w_{s}$, respectively, as well as the fire spread rate due to fuel type, denoted by $r_{f}$. These parameters are stochastic and can vary throughout the environment. Therefore, each vertex may be affected differently depending on the fire front’s extent and the environmental variability. Specifically, we denote $\theta(i)\in[0,2\pi]$ as the wind direction affecting vertex $i$, the wind speed at the location of vertex $i$ as $w_{s}(i)\in\mathbb{R}^{+}$, and the fire spread rate as $r_{f}(i)\in\mathbb{R}^{+}$. Here, $\theta(i),w_{s}(i),r_{f}(i)$, $\forall i$, are random realizations of the wind direction, wind speed, and fire spread rate at the location of vertex $i$. The parameters $\alpha_{1}(i)$, $\alpha_{2}(i)$, and $\alpha_{3}(i)$ denote the shape parameters governing the elliptical fire growth from vertex $i$, representing respectively the lengths of the semi-minor axis, the semi-major axis, and the distance from the ignition point to the center of the ellipse, defined respectively as: $\alpha_{1}(i)=\frac{r_{f}(i)+\frac{r_{f}(i)}{HB(i)}}{2LB(i)}$, $\alpha_{2}(i)=\frac{r_{f}(i)+\frac{r_{f}(i)}{HB(i)}}{2}$, and $\alpha_{3}(i)=\alpha_{2}(i)-\frac{r_{f}(i)}{HB(i)}$, where $HB(i)$ is the head-to-back ratio accounting for the difference between the fire’s forward (head) and backward (back) spread from the ignition point, while $LB(i)$ is the length-to-breadth ratio which determines the overall elongation of the fire’s elliptical shape. These are defined as: $HB(i)=\frac{LB(i)+(LB(i)^{2}-1)^{0.5}}{LB(i)-(LB(i)^{2}-1)^{0.5}}$, and $LB(i)=0.936\exp(0.2566w_{s}(i))+0.461\exp(-0.1548w_{s}(i))-0.397$. For a more in-depth description these parameters we refer the reader to [4]. Subsequently, the propagation of the fire front process $X_{t}=[x^{1}_{t},\ldots,x^{N}_{t}]^{\top}\in\mathcal{X}$ is more compactly expressed as: $X_{t}=\xi(X_{t-1},E_{t-1})$, where $E_{t}\sim P_{E}$, with $E_{t}\in[0,2\pi]^{N}\times[0,\infty)^{N}\times[0,\infty)^{N}$, denotes a random realization of $\{\theta(i),w_{s}(i),r_{f}(i)\}_{i=1}^{N}$ drawn from the PDF $P_{E}$, which captures the stochasticity of the fire front propagation acting as a stationary process noise.
II-C Agent Dynamics and Sensing Model
An autonomous mobile agent represented by a point-mass object, evolves inside a bounded planar environment $\mathcal{E}\subset\mathbb{R}^{2}$ according to discrete-time dynamics of the form $y_{t}=f_{a}(y_{t-1},u_{t-1})$ [10], where $y_{t}\in\mathcal{Y}$ is the state of agent at time $t$, and $u_{t}\in\mathcal{U}$ is the control input. In addition, the agent has a finite sensing range for observing its surroundings (i.e., through a camera), which is given by a circular region with radius $R_{a}$ i.e., $O_{t}=\{x\in\mathbb{R}^{2}\mid\|x-y^{p}_{t}\|\leq R_{a}\}$, where $y^{p}_{t}$ is the agent’s position at time $t$.
The agent uses its camera to observe the state of the fire front, i.e., by taking snapshots and determining the location of the fire front from the image snapshots using image processing (i.e., object detection). Due to sensing and image processing imperfections, this process carries a certain degree of inaccuracy, resulting in noisy observations. Specifically, for fire front vertex $i$ with true state $x^{i}_{t}$, the agent observes the measurement $z^{i}_{t}$ inside its sensing range according to: $z^{i}_{t}=h(x^{i}_{t})+w^{i}_{t},$ where $h(\cdot)$ is a function that relates the true states to the received measurements, and $w^{i}_{t}\sim\mathcal{N}(0,\sigma^{2}_{z}I_{2\times 2})$ represents measurement noise. The noise is independent and identically distributed (i.i.d.) according to a zero-mean Gaussian distribution with variance $\sigma^{2}_{z}$, where $I_{2\times 2}$ is the $2\times 2$ identity matrix. Additionally, $w^{i}_{t}$ is independent of the process noise described in the previous section.
The object detection algorithm often produces multiple fragmented pixel blobs for the same fire-front vertex, so that $x^{i}_{t}$ is associated with the measurement vector $z^{i,1}_{t},\ldots,z^{i,n^{i}_{t}}_{t}]$. The number of such blobs depends on the true vertex location, and is modeled as a Poisson random variable with rate $\lambda^{i}_{t}(x^{i}_{t})$. Thus, a vertex $x^{i}_{t}$ may yield $n^{i}_{t}\sim\text{Pois}(\lambda^{i}_{t}(x^{i}_{t}))$ detections within the sensing range. The resulting measurement set follows a Poisson point process [14] with intensity
$\gamma^{i}_{t}(z|x^{i}_{t},y_{t})=\lambda^{i}_{t}(x^{i}_{t})\,p(z|x^{i}_{t}),\quad x^{i}_{t}\in O_{t},$ (3)
and $\gamma^{i}_{t}(z|x^{i}_{t},y_{t})=0$ otherwise, where $p(z|x^{i}_{t})=\mathcal{N}(z;h(x^{i}_{t}),\sigma^{2}_{z}I_{2\times 2})$ is the normalized measurement likelihood restricted to $O_{t}$.
III Adaptive Monitoring via Information-seeking Predictive Control
III-A Fire Front Recursive State Estimation
Bayesian recursive state estimation for systems with fixed-dimensional state and measurement vectors is formulated through the predictor-corrector recursion:
$\mathcal{B}^{-}_{t}(x_{t}|z_{1:t-1})=\int p(x_{t}|x_{t-1})\,\mathcal{B}_{t-1}(x_{t-1}|z_{1:t-1})\,dx_{t-1},$
$\mathcal{B}_{t}(x_{t}|z_{1:t})=\frac{p(z_{t}|x_{t})\,\mathcal{B}^{-}_{t}(x_{t}|z_{1:t-1})}{\int p(z_{t}|x_{t})\,\mathcal{B}^{-}_{t}(x_{t}|z_{1:t-1})\,dx_{t}},$ (4)
where with slight abuse of notation in Eq. (III-A) $x_{t}\in\mathbb{R}^{d_{x}}$ is the state of the system, $z_{t}\in\mathbb{R}^{d_{z}}$ is the received measurement, $p(x_{t}|x_{t-1})$ is the transitional density governed by the stochastic process dynamics, $p(z_{t}|x_{t})$ is the measurement likelihood function, $\mathcal{B}^{-}_{t}(x_{t}|z_{1:t-1})$ is the predictive belief distribution at time $t$, and $\mathcal{B}_{t}(x_{t}|z_{1:t})$ is the posterior belief of $x_{t}$ when all measurements $z_{1:t}=[z_{1},\ldots,z_{t}]$ up to time $t$ have been received. Subsequently, given the recursion in Eq. (III-A) the minimum mean square estimator (MMSE) $\hat{x}^{\text{MMSE}}_{t}$ is given by: $\hat{x}^{\text{MMSE}}_{t}=\int x_{t}\,\mathcal{B}_{t}(x_{t}|z_{1:t})\,dx_{t}$.
However, in our problem, at time $t$ the observation is a point pattern of random cardinality, not a fixed-length vector, hence the standard vector-likelihood underlying Eq. (III-A) is not directly applicable. We therefore replace the likelihood in Eq. (III-A) with the appropriate set-likelihood and apply Bayes’ rule with that form. To achieve this, first observe that the transitional density $p(x_{t}|x_{t-1})$ in Eq. (III-A) becomes $p(X_{t}|X_{t-1})=\int\delta(X_{t}-\xi(X_{t-1},E_{t-1}))\,P_{E}(E_{t-1})\,dE_{t-1}$ as direct consequence of the fire front stochastic dynamics, where $\delta(\cdot)$ is the Dirac delta function. Subsequently, we can compute the predicted belief $\mathcal{B}^{-}_{t}(X_{t}|Z_{1:t-1})$ at time step $t$, assuming the posterior density at the previous time step, $\mathcal{B}_{t-1}(X_{t-1}|Z_{1:t-1})$, is known. The posterior belief $\mathcal{B}_{t}(X_{t}|Z_{1:t})$ can then be obtained by incorporating the measurement set $Z_{t}$ via the correction step shown in Eq. (III-A), provided that an expression for the measurement likelihood function $p(Z_{t}|X_{t},y_{t})$ is available. This process can then be recursively applied to the next time step.
Proposition: Let $X_{t}=[x^{1}_{t},\ldots,x^{N}_{t}]^{\top}$ be the state of the fire front at time $t$, and let $Z_{t}=[z^{1}_{t},\ldots,z^{m_{t}}_{t}]$ denote the received measurement vector at time step $t$. The likelihood of $Z_{t}$ given $X_{t}$ and $y_{t}$ is given by: $p(Z_{t}|X_{t},y_{t})=$
$\frac{1}{m_{t}!}\exp\left(-\sum_{i=1}^{N}\lambda_{i}(x^{i}_{t})\right)\prod_{k=1}^{m_{t}}\left(\,\sum_{i=1}^{N}\gamma^{i}_{t}\bigl(z_{k}|x^{i}_{t},y_{t}\bigr)\right),$ (5)
where $\lambda^{i}_{t}(x^{i}_{t})=\int_{O_{t}}\gamma^{i}_{t}(z\,|\,x^{i}_{t},y_{t})\,dz$ is the expected number of detections from vertex $i$, $\gamma^{i}_{t}(z\,|\,x^{i}_{t},y_{t})$ is the Poisson intensity corresponding to vertex $i$ within the sensing region $O_{t}$ as defined in Eq. (3), and $m_{t}!$ denotes the factorial of $m_{t}$. Consequently, $p(Z_{t}|X_{t},y_{t})$ can be used directly in Eq. (III-A) enabling the handling of multiple objects and multiple measurements without requiring explicit measurement-to-object association, and allowing the computation of the MMSE as discussed previously.
III-B Information-seeking Predictive Control
The problem in Sec. II-A is addressed via the information-seeking predictive controller shown in Problem (P1), formulated as a receding horizon SOC problem. The goal is to compute control inputs $\{u_{t+\tau-1|t}\}_{\tau=1}^{T}$ that optimize the agent’s sensing behavior by minimizing the cumulative uncertainty in the (pseudo) posterior beliefs of the fire front states, as defined in Eq. (6a). At each time step $t$, only the first control input $u_{t|t}$ is applied, and the process is repeated over a shifted horizon.
For each prediction time step $\tau\in\{1,\ldots,T\}$ in the horizon, the agent predicts the fire front state $X_{\tau}$ forward in time using the Bayesian prediction step, based on the transition density $p(X_{\tau}|X_{\tau-1})$ and the (pseudo) posterior belief from the previous time step, $\mathcal{B}_{t+\tau-1|t}(X_{\tau-1}|Z_{1:t+\tau-1|t})$, as shown in Eqs. (6b)-(6c). The constraints in Eqs. (6d)-(6e) arise from the agent dynamics, which are assumed to be deterministic in this work. At the predicted time step $\tau$, given the agent’s predicted state $y_{t+\tau|t}$ and sensing range $O_{t+\tau|t}$, the agent receives the predicted measurement set $Z_{t+\tau|t}$. This set is then used to compute the posterior belief $\mathcal{B}_{t+\tau|t}(X_{\tau}|Z_{1:t+\tau|t})$ via the Bayesian correction step, as shown in Eq. (6f), using the joint likelihood function $p(Z_{t+\tau|t}|X_{\tau},y_{t+\tau|t})$ and the predicted density. This posterior distribution subsequently becomes the prior for the next prediction step, continuing the recursive process. The predicted measurements $Z_{t+\tau|t}$ represent hypothetical observations based on the planned control inputs and the anticipated fire front state. Since actual measurements are only available after executing the control actions, the objective in Eq. (6a) requires taking an expectation over all possible future measurement sequences. This enables informed decision-making by accounting for potential outcomes without executing the corresponding trajectories.
Problem (P1):Information-seeking Predictive Control
$\hskip-8.53581pt\underset{\{u_{t+\tau-1|t}\}_{\tau=1}^{T}}{\min}\mathbb{E}_{Z_{t}^{1:T}}\left\{\sum_{\tau=1}^{T}\nu^{\tau}\mathcal{C}_{\tau}\!\left(\mathcal{B}_{t+\tau|t}(\cdot|Z_{1:t+\tau|t})\right)\right\}$ (6a)
subject to:
$\hskip-8.53581pt\mathcal{B}^{-}_{t+\tau|t}=\int p(X_{\tau}|X_{\tau-1})\,\mathcal{B}_{t+\tau-1|t}(X_{\tau-1}|\cdot)\,dX_{\tau-1},$ (6b)
$\hskip-8.53581pt\mathcal{B}_{t|t}=\mathcal{B}_{t|t-1},$ (6c)
$\hskip-8.53581pty_{t+\tau|t}=f_{a}(y_{t+\tau-1|t},u_{t+\tau-1|t}),$ (6d)
$\hskip-8.53581pty_{t|t}=y_{t|t-1},$ (6e)
$\hskip-8.53581pt\mathcal{B}_{t+\tau|t}(\cdot|Z_{1:t+\tau|t})\propto p\left(Z_{t+\tau|t}|X_{\tau},y_{t+\tau|t}\right)\mathcal{B}^{-}_{t+\tau|t},$ (6f)
$\hskip-8.53581pty_{t}\in\mathcal{Y},u_{t}\in\mathcal{U},X_{t}\in\mathcal{X},Z_{t}\in\mathcal{Z},$ (6g)
$\hskip-8.53581pt\mathcal{C}_{\tau}\in[0,1],\nu\in(0,1],\tau=\{1,\ldots,T\}.$ (6h)
Equations (6b)-(6f) admit no closed form: the model is nonlinear and non-Gaussian with set-valued (multi-object, multi-measurement) observations, so Kalman-type filters are inapplicable. Therefore, the recursion is implemented using Sequential Importance Resampling (SIR), i.e., particle filtering. Specifically, the belief $\mathcal{B}_{\tau}$ is represented by a set of weighted particles $\mathcal{B}_{\tau}=\{w^{(i)}_{\tau},X^{(i)}_{\tau}\}_{i=1}^{N_{s}}$, where $X^{(i)}_{\tau}=[x^{1}_{\tau},\ldots,x^{N}_{\tau}]^{\top}$. These particles are propagated to the next time step according to the process dynamics and reweighted using the likelihood function to compute the posterior. For notational convenience, we will often write $t+\tau|t$ as $\tau$ when no ambiguity arises. The functional $\mathcal{C}_{t+\tau|t}:\mathcal{B}_{t+\tau|t}(X_{t+\tau|t}|Z_{1:t+\tau|t})\rightarrow[0,1]$ in Eq. (6a) quantifies the uncertainty of the fire-front state $X_{t+\tau|t}$ encoded in the posterior distribution at time $t+\tau|t$ , conditioned on all hypothetical measurements $Z_{1:t+\tau|t}$. This uncertainty is measured by the Risk-Weighted Dispersion (RWD) defined as:
$\mathcal{C}_{t+\tau|t}(\mathcal{B}_{t+\tau|t})=\frac{1}{\omega}\sum_{\varepsilon\in\tilde{\mathcal{E}}}\mathcal{R}(\varepsilon)\det(\Sigma^{\varepsilon}_{t+\tau|t}),$ (7)
where the environment $\mathcal{E}$ is discretized in space to form a 2D grid $\tilde{\mathcal{E}}$, composed of a finite number of non-overlapping, equally sized cells $\tilde{\mathcal{E}}=\{\varepsilon_{1},\ldots,\varepsilon_{|\tilde{\mathcal{E}}|}\}$, such that $\bigcup_{i=1}^{|\tilde{\mathcal{E}}|}\varepsilon_{i}=\mathcal{E}$. The term $\mathcal{R}(\varepsilon)\in[0,1]$ denotes the risk value associated with cell $\varepsilon$, reflecting the severity of fire presence in that region. The random quantity $\det(\Sigma_{t+\tau|t}^{\varepsilon})$ is the determinant of the sample covariance matrix $\Sigma_{t+\tau|t}^{\varepsilon}$ computed from all particle points residing in cell $\varepsilon$ at time step $t+\tau|t$, and $\omega$ is a scaling factor ensuring that $\mathcal{C}_{t+\tau|t}(\mathcal{B}_{t+\tau|t})\in[0,1]$. Finally, the parameter $\nu\in(0,1]$ in Eq. (6a) is a discount factor that controls the relative importance of future decisions.
III-C Adaptive LCB-guided Policy Search
Problem (P1) is a stochastic, multi-dimensional, non-linear, and non-convex optimization problem that cannot be directly solved in its original form. However, we can address an equivalent version by reformulating (P1) as a Markov Decision Process (MDP) [16]. To achieve this, we assume that the agent’s control inputs $u_{t}\in\mathcal{U}$ can be reduced to a finite set $\mathbb{U}$, consisting of $|\mathbb{U}|$ discrete control vectors $\hat{u}_{t}\in\mathbb{U}$. This discretization, in turn, leads to a finite set of possible agent states $\hat{y}_{t}\in\mathbb{Y}_{t}$. Consequently, (P1) can be reformulated as an MDP $\langle\mathcal{S},\mathbb{U},\mathcal{T},\mathcal{C}\rangle$, where $\mathcal{S}$ is the state space of the system, and an individual state $s\in\mathcal{S}$ is represented as the tuple $s_{t}=(\mathcal{B}_{t},\hat{y}_{t})$. Note that the fire front process evolves independently of the agent’s actions. The transition function $\mathcal{T}:\mathcal{S}\times\mathbb{U}\rightarrow\mathcal{S}$ describes the evolution of the system in response to agent actions. Although the agent’s actions are deterministic in our setting, the transition function $\mathcal{T}$ remains stochastic due to the randomness in the agent’s observations upon executing an action. Specifically, we have: $\mathcal{T}:p\left(s^{\prime}_{t}=(\mathcal{B}^{\prime},\hat{y}^{\prime})\mid s_{t-1}=(\mathcal{B},\hat{y}),\hat{u}_{t-1}=u\right)$. The cost function $\mathcal{C}$ assigns a cost to a specific state $s^{\prime}$ resulting from applying action $\hat{u}$ at state $s$. With slight abuse of notation, we denote it as $)\mathcal{C}(s^{\prime}=(\mathcal{B},\hat{y}$)), which effectively operates on the posterior belief $\mathcal{B}$ in state $s^{\prime}$, derived from the agent state $\hat{y}$. The cost is defined according to Eq. (7).
We define a finite-horizon open-loop policy over $T$ steps as the control input sequence $\pi=\{\hat{u}_{t+\tau-1|t}\}_{\tau=1}^{T}\in\mathbb{U}^{T}$. Let $\Pi_{T}=\mathbb{U}^{T}$ denote the set of all such admissible policies. Each policy is a predetermined sequence of actions evaluated via simulation of the MDP’s stochastic state transitions. Given that the agent starts from an initial state $s_{t}=(\mathcal{B}_{t},\hat{y}_{t})$ at time step $t$, a policy $\pi\in\Pi_{T}$ can then be simulated under the MDP to obtain:
$V^{\pi}_{t}(s_{t})=\mathbb{E}\left\{\sum_{\tau=1}^{T}\nu^{\tau}\mathcal{C}_{\tau}\left(s^{\pi}_{t+\tau|t}\right)\right\},$ (8)
where $s^{\pi}_{t+\tau|t}$ is the state encountered at level $\tau$ in the horizon under policy $\pi$ when starting from state $s_{t}$. Subsequently, the optimal policy that minimizes the expected cumulative cost over the horizon is given by $\pi^{\star}=\arg\min_{\pi\in\Pi_{T}}\,V^{\pi}_{t}(s_{t})$. To solve this problem we build upon the Upper Confidence Bound 1 (UCB1) framework [2] and we treat each sequence $\pi=\{\hat{u}_{t+\tau-1|t}\}_{\tau=1}^{T}\in\mathbb{U}^{T}$ as an arm in a multi-armed bandit. We use rollouts to simulate policy outcomes and employ a UCB1-like adaptive selection strategy to efficiently balance exploration and exploitation.
Algorithm 1 Adaptive LCB-guided Search
1:Input: Policy set $\Pi_{T}$, Initial state $s_{t}=(\mathcal{B}_{t},\hat{y}_{t})$
2:Initialize $I^{0}(\pi)=0,\forall\pi\in\Pi_{T}$
3:for $n=1,\dots,(n_{\max}\geq|\Pi_{T}|)$ do
4:  Sample policy: $\hat{\pi}\!=\!\arg\min\mathrm{LCB}^{n}(\pi,s_{t})$ (Eq. (9))
5:  for $\tau=1,\dots,T$ do
6:   Compute predictive density: $\mathcal{B}^{-}_{t+\tau|t}$ via Eq. (6b)
7:   Move to new state: $\hat{y}_{t+\tau|t}=f_{a}\left(\hat{y}_{t+\tau-1|t},\hat{\pi}_{\tau}\right)$
8:   Sample meas.: $Z_{t+\tau|t}\sim p(Z|X_{t+\tau|t},\hat{y}_{t+\tau|t})$
9:   Compute posterior: $\mathcal{B}_{t+\tau|t}$ via Eq. (6f)
10:   Compute stage cost: $c_{\tau}=c_{\tau-1}+\nu^{\tau}\mathcal{C}_{\tau}(\mathcal{B}_{t+\tau|t})$
11:  end for
12:  Update LCB: $Q^{n}(\hat{\pi},s_{t})=\mathrm{mean}\left(Q^{n-1}(\hat{\pi},s_{t}),c_{T}\right)$
13:                           $I^{n}(\hat{\pi})=I^{n}(\hat{\pi})+1$
14:end for
15:Output: Optimal policy $\pi^{\star}=\arg\min~\mathrm{LCB}^{n_{\text{max}}}(\pi,s_{t})$
Figure 1: Simulation Setup: (a) Fire front true evolution, (b)(c) Wind speed parameters, (d) Risk map, and (e)(f) Fire spread rate parameters.
The proposed adaptive-search algorithm iteratively computes a Lower Confidence Bound (LCB) on the expected total cost associated with each $T$-finite control sequence, which is then used to adaptively select the next policy to simulate. Specifically, for an agent at state $s_{t}$, we define the LCB score of a policy $\pi\in\Pi_{T}$ at iteration $n$ as:
$\mathrm{LCB}^{n}(\pi,s_{t})=\begin{cases}Q^{n}(\pi,s_{t})-\sqrt{\frac{2\ln n}{I^{n}(\pi)}},&\text{if }I^{n}(\pi)\neq 0,\\ \mathrm{LCB}^{n}_{\min},&\text{otherwise},\end{cases}$ (9)
where $I^{n}(\pi)$ denotes the number of times policy $\pi$ has been simulated up to iteration $n$, and $Q^{n}(\pi,s_{t})=\frac{1}{I^{n}(\pi)}\sum_{i=1}^{n}\delta(\pi^{i}-\pi)L(\pi,s_{t})$ is the sample mean of the cumulative cost incurred by policy $\pi$. Here, $L(\pi,s_{t})=\sum_{\tau=1}^{T}\nu^{\tau}\mathcal{C}_{\tau}\left(s^{\pi}_{t+\tau|t}\right)$ is the total discounted cost of executing policy $\pi$ starting from state $s_{t}$, and $\delta(\cdot)$ is the discrete Dirac delta function. The constant $\mathrm{LCB}^{n}_{\min}<-\sqrt{2\ln n}$ ensures that any policy not yet selected by iteration $n$ will have an LCB value lower than that of any previously selected policy, thereby guaranteeing it will be explored with higher probability in subsequent iterations. Finally, the term $\sqrt{\frac{2\ln n}{I^{n}(\pi)}}$ serves as an exploration bonus, encouraging the selection of under-explored policies. Under the standard UCB1 assumptions, the expected fraction of iterations on which LCB selects a suboptimal policy is $O\!\big(\tfrac{\ln n}{n}\big)$; hence the optimal policy is selected with asymptotic frequency $1$ in expectation as $n\to\infty$.
The complete LCB-guided adaptive search algorithm is shown in Algorithm 1. At each time step $t$, the algorithm identifies the optimal policy $\pi^{\star}$ over the horizon $\{t+\tau|t\}_{\tau=1}^{T}$. The agent then executes the first control input of $\pi^{\star}$, transitions to a new state, receives the “real” measurement, computes the posterior belief $\mathcal{B}_{t+1}$, and the algorithm is re-applied at the next time step $t+1$. We should note that intelligent pruning techniques, such as $\epsilon$-suboptimal reductions [1], can be designed and integrated into the proposed approach to focus on the most promising set of control inputs at each time step thereby reducing the runtime complexity.
IV Evaluation
IV-1 Simulation Setup
[…]
Figure 2: Performance Evaluation: (a) Agent trajectory ($\star$ and $\times$ indicate start and stop states respectively), and fire front state (i.e., estimated with solid line, and true state with dotted line) obtained with Alg. 1, over 25-steps experiment, with myopic settings (i.e., horizon $T=1$), (b) Same result with non-myopic settings (i.e. horizon $T=3$), (c) Estimation error for different configurations of the proposed approach and comparisons with baselines.
IV-2 Results
Figure 1(a) illustrates the true evolution of the fire front over a simulation period of $T_{s}=25$ time steps, influenced by the environmental conditions described above. The fire-spread physics in this model support propagation in all directions from the ignition point. As a result, even under strong wind and fuel conditions, the upwind (back) perimeter of the fire continues to advance, albeit at a slower rate, as depicted. Furthermore, Fig. 1(a)-(c) illustrates how variations in wind direction, wind speed, and spread rate influence the propagation of the fire front, either by accelerating or decelerating its advancement.
The output of Alg. 1 for this setup is shown in Fig. 2(a) and Fig. 2(b), corresponding to horizon lengths of $T=1$ (myopic) and $T=3$ (non-myopic), respectively. The agent is initialized at position $(x,y)=(700,800)$, as indicated by the green asterisk, running Alg. 1 in a rolling horizon fashion, as discussed in Sec.III-C. The green line in the figures is the agent’s final trajectory over 25 time steps, resulting from the minimization of the expected cumulative RWD over the planning horizon at each time step. The dotted black lines show the true evolution of the fire front, while the time color-coded lines represent the estimated front; ideally, these two should align closely. As shown, the non-myopic approach plans multiple steps ahead and achieves improved performance compared to the myopic strategy, which plans greedily. Specifically, the myopic behavior of the agent in Fig. 2(a) prevents it from targeting the high-risk region in the top-left corner of Fig. 1(d), which in this scenario is also associated with high uncertainty, as illustrated in Figs. 1(c) and 1(f). This limitation results in significant estimation errors, as shown.
Subsequently, Fig. 2(c) illustrates the performance of the proposed approach in terms of root mean square error (RMSE) i.e., $\sqrt{\mathbb{E}\big((\hat{X}_{t}-X_{t})^{2}\big)}$, between the estimated fire front state $\hat{X}_{t}$ and the true state $X_{t}$ over 25 time steps, using a uniform risk map. Specifically, we conducted 50 Monte Carlo trials, randomly initializing both the fire front and the agent’s position within the simulation environment described earlier. The figure presents the average $\log_{10}\mathrm{RMSE}$ per time step per vertex over the 25-step simulation, comparing six different approaches. The baseline, denoted as Inf. SR, corresponds to an agent with infinite sensing range that remains stationary and simply runs the particle filter. In this case, the RMSE arises solely from measurement noise and multiple detections, with no influence from control actions-representing the best achievable performance under the given settings. The Static approach involves a stationary agent with a finite sensing range, while the Random approach uses an agent also with finite sensing range that selects a random control input at each time step. As expected, both approaches result in significant errors. The figure also includes the proposed method evaluated with different planning horizon lengths, i.e., $T=1$ (myopic), and non-myopic settings with $T=3$ and $T=5$, and clearly demonstrates improved performance as the planning horizon increases. Note that the baseline is unattainable in this setting due to the agent’s limited sensing capabilities.
V Conclusion
This paper considers the problem of fire front monitoring under uncertainty by formulating it as a stochastic optimal control problem that integrates sensing, estimation, and control. A recursive Bayesian estimator was developed for elliptical-growth fire front processes, and the control problem was formulated as a finite-horizon Markov Decision Process (MDP). An information-seeking control law was then derived using a lower confidence bound (LCB)-based adaptive search, enabling optimal risk-aware planning.